using the five-step problem-solving process
A rectangular support unit in a television is made with dimensions measuring 6 centimeters by 7 centimeters. A new model being designed for the upcoming year will have a larger rectangular support with dimensions measuring 6.3 centimeters by 7.4 centimeters. By how much will the area of the rectangular support be increased?the formula i was using was w*l but im not getting 5 steps can some one help me

Answer by mananth(16946)

original support is 7 by 6 cm
so area = 42 cm^2
New dimensions 6.3 by 7.4 cm
area = 6.3*7.4 =46.62cm^2
The difference is the increase in area
46.62-42=4.62cm^2